OpenStudy (anonymous):

Given the following functions f(x) and g(x), solve f over g (−5) and select the correct answer below: f(x) = 2x − 20 g(x) = x − 1 −5 5 one sixth 30

myininaya (myininaya):

\[(\frac{f}{g})(5)=\frac{f(5)}{g(5)}\]

myininaya (myininaya):

f(5) can be found by using f(x)=2x-20 and replacing x with 5

myininaya (myininaya):

g(5) can be found by using g(x)=x-1 and replacing x with 5

myininaya (myininaya):

then just put what you got for f(5) over what you got for g(5)

myininaya (myininaya):

oh didn't see 5 was -5 actually replace my 5's with -5's

OpenStudy (anonymous):

would that make it positive ?

OpenStudy (anonymous):

b?

myininaya (myininaya):

so what did you get for f(-5) that is 2(-5)-20? and what did you get for g(-5) that is -5-1?

OpenStudy (anonymous):

-30 -6

myininaya (myininaya):

cool and -30/-6 is a positive number that positive number is yes 5 good job

OpenStudy (anonymous):

thank you

myininaya (myininaya):

np! you did all the work! :)